(*) Find the resistance in ohms of an aluminum wire that is 15 m long and has diameter 0.0025 m.(**) suppose the current in the wire causes the temperature of the wire to increase 20 degree C. what is the new resistance of the wire?(p = 2.82 x 10^-8 ohms x m, a = 3.9 x 10^-3/degree C)
Answer #1

r = 0.0025/2 m = 1.25*10^-3 m
L = 15 m
A =pi*r^2
= pi* (1.25*10^-3)^2
=4.91*10^-6 m^2

R = rho*L/A
= (2.82*10^-8)*15/(4.91*10^-6)
= 0.086 ohm
Answer: 0.086 ohm

----------------
use:L
R' = R* [1+a*delta T]
= 0.086 * [1 + 3.9*10^-3* (20)]
= 1.81 ohm
Answer: 1.81 ohm